# FoMEMO: Towards Foundation Models for Expensive Multi-objective Optimization

### v1

@@ 5.2 Parallel Scalability
In addition to sequential optimization, we further investigate the scalability of our proposed method in the context of parallel optimization. We demonstrate the parallel optimization performance of all algorithms when generating $q=5$ candidate solutions in each iteration as an example, as shown in Figure 3. More results can be found in the appendix. Excluding the initial sample point, a total of 20 iterations are conducted. For preference-based acquisition functions, we iterate over $q$ preferences to generate $q$ candidate solutions in each iteration. For preference-free methods, using a limited number of preferences to approximate the hypervolume improvement introduces a degree of deviation. Therefore, we repeatedly perform the UHVI calculation $q$ times, resulting in $q$ candidates for parallel evaluation. Our findings indicate that, despite employing a straightforward parallelization strategy, our algorithms remain competitive when compared to existing state-of-art parallelization methods. More principled parallel acquisition functions and sample selection strategies deserve further exploration in the future.
5.3 Runtime Advantage
Figure 4: Comparison of query time of different model-based algorithms in sequential optimization. We record the average total time in seconds (using log scaling) taken by different algorithms to generate candidate solutions under 10 independent runs.
We compare the query time of different algorithms in sequential optimization across all real-world applications, as depicted in Figure 4. The query time encompasses only the sum of duration required to optimize the acquisition functions based on the models when generating candidate solutions, excluding the time spent on model updates (if applicable) and the time dedicated to evaluating the true functions. The comparisons are conducted on the same device, and the reported results are averages from 10 independent runs. The experimental results demonstrate the superiority of the unique in-context optimization of our methods in terms of the query time. Overall, the inclusion of expensive hypervolume calculations leads to the indicator-based algorithms qEHVI and qNEHVI experiencing the longest query time, especially for problems with higher objective dimensions. In contrast, benefiting from the fast inference of the foundation model, the in-context optimization of all acquisition functions of FoMEMO demonstrates orders of magnitude efficiency. Among these, FoMEMO-UHVI requires the least amount of time, as it samples fewer preferences to efficiently approximate the hypervolume indicator. However, this efficiency comes at the cost of some optimization performance. In contrast, FoMEMO-UCB has achieved a favorable balance between optimizing performance and running time.
@@ 6 Conclusion and Future Work
We propose a novel paradigm named FoMEMO, which, to the best of our knowledge, represents the first attempt to establish a foundation model to address a wide range of expensive MOPs in an in-context manner. Extensive experiments verify that our method exhibits superior adaptability and generalization in both synthetic problems and real-world applications, showing competitive performance in most cases compared to existing methods.
@@
In the future, we expect that this work opens up opportunities for many exciting directions worth exploring: (i) Develop architectures that allow for datasets with larger sizes, including embedding techniques and training schemes that efficiently handle higher-dimensional feature and objective spaces. (ii) Develop efficient acquisition functions that adapt to high-dimensional and parallel settings. (iii) Explore alternative methods for generating synthetic data, such as integrating efficient calculations of hypervolume or R2 (Hansen and Jaszkiewicz 1994; Brockhoff, Wagner, and Trautmann 2012) indicators as label responses to support efficient training. (iv) Investigate whether fine-tuning can help improve the performance in domain-specific tasks.
Appendix
Appendix A Limitations and Potential Future Work
A.1 GP Priors Limitations
In this paper, we focus on addressing a general class of unconstrained multi-objective expensive optimization problems characterized by continuous, non-linear objective functions that exhibit smoothness and can incorporate uncertainty, making them well-suited for Gaussian Process (GP) modeling. Therefore, we can extensively sample from GP priors to generate a substantial amount of synthetic training data, simulating a wide range of potential mappings that may exist in such real-world scenarios. The experimental results further demonstrate the applicability and generalization of our method. However, when confronted with problems that involve special knowledge completely beyond GP priors, our method may struggle to generalize to these specific types of cases. Potential solutions include incorporating expert knowledge to design specialized priors during synthetic data generation, or interacting with the evaluation environment during optimization to discover and enrich potential priors. Since synthetic data can be generated from the designed priors through extensive sampling, these enhancements can be seamlessly integrated into our existing method. Furthermore, instead of permanently freezing the parameters after a single round of synthetic pre-training, it may be advantageous to fine-tune the pre-trained model using a small amount of domain-specific data.
A.2 Scalability Challenges
Although we have evaluated the scalability of our approach across problems with varying feature and objective spaces (see Section E.3), those involving significantly higher-dimensional features or objectives pose considerable challenges. This scalability issue is also widely recognized as one of the main difficulties within the multi-objective optimization community. To this end, several promising avenues worth further exploration. In the context of high-dimensional feature spaces, the development of efficient dimensionality reduction methods and embedding techniques, along with effective search strategies such as the trust region Bayesian optimization (TuRBO) algorithm (Eriksson et al. 2019), presents significant potential. For problems involving a large number of objectives, the number of solutions required to approximate the entire Pareto set increases exponentially, resulting in significant overhead for both optimization and decision-making. Therefore, rather than pursuing a comprehensive set of Pareto solutions, recent work has focused on identifying a small set of representative solutions, ensuring that each objective can be well addressed by at least one solution within the set (Lin et al. 2024; Liu et al. 2024). This motivates the development of foundational models and parallel acquisition functions based on the set scalarizations (Lin et al. 2024).
Appendix B Potential Societal Impact
The proposed method may yield the following positive impacts:
  • It provides a general paradigm that can quickly adapt to numerous expensive multi-objective challenges continuously emerging in the real world, enabling users to efficiently identify superior solutions in a sample-efficient manner while minimizing the expenditure of social resources.
  • The generation of the training corpus does not require access to any real-world data, showcasing significant potential for a variety of real-world scenarios where costly experimental data is severely limited or where user privacy and commercial copyright issues are critical.
  • Our approach offers a user-friendly and flexible interactive platform. Users can receive uncertainty-guided feedback on any sub-problem according to their individual preferences, allowing them to explore and exploit customized solutions that better align with their needs.
On the negative side, we acknowledge that we have invested considerable computational resources in implementing the algorithm framework and tuning the hyperparameters of the pre-trained network. However, we believe that our work will significantly lower the foreseeable design barriers and computational costs for practitioners in the future, which is one of the primary goals of our work. We hope that the societal benefits across various fields from directly adapting our method to few-shot scenarios, or even fine-tuning it with limited domain knowledge, will far outweigh the costs incurred in our efforts.
Appendix C Model and Algorithm Details
C.1 Details of Foundation Model and Pre-training
@@
Maximum Sample Length 128
Model Training Batch Size 256
@@
Learning Rate 0.0001
Our foundational model mainly employs the recently proposed Prior-Data Fitted Networks (PFNs) (Müller et al. 2021) as the basic architecture, which is based on a Transformer encoder (Vaswani et al. 2017). We utilize linear layers to encode trajectory and query points, along with sampled preferences, as contextual inputs for the foundation model. Following the methods outlined in (Hollmann et al. 2022) and (Müller et al. 2023), we set a maximum dimension $K$ for encoding the inputs, which are applicable to both feature and objective dimensions. We zero-pad the sampled dimension $k$ and linearly scale the dimension by a factor of $K/k$ to ensure that the magnitudes of the input encodings are consistent across different dimensions. Since different query points should not depend on one another, an attention mask is implemented to allow each trajectory point can only attend to each other, while the query points can only attend to the trajectory points. To maintain the permutation invariance of the input dataset, the positional encodings are removed and the input points are fed to the model as the sum of their encoded embeddings. During training, we maintain a maximum sample length of $N$ with $n$ trajectory points and NnN-n query points, as used in (Müller et al. 2021). We sample each $n$ from $\{1,...,N-1\}$ with the weight of 1/(Nn)1/(N-n) to simulate varying trajectory sizes during the optimization process, while ensuring that each trajectory size has approximately the same number of query points for training. To model the continuous distribution for prediction, the regression head of PFNs discretizes the continuous distribution into a piecewise constant distribution, referred to as the Riemann distribution (Müller et al. 2021). This helps transform the model training from a regression task to a classification problem. The boundaries of each discrete interval are estimated by sampling a large synthetic dataset from the same synthetic prior, ensuring that each interval maintains equal probability in the prior data. Additionally, the last bar on each side is replaced with a suitably scaled half-normal distribution to enhance training stability.
@@
where $\sigma$ and $l$ represent the output scale and length scale kernel parameters, respectively. We also sample the length scale parameter for each feature dimension during the latent function sampling, along with an random observation noise to enhance function diversity. In the third step, we randomly sample input features from the $d$-dimensional unit cube, and then evaluate the sampled latent functions to generate the corresponding observations, resulting in $n$ trajectory samples and NnN-n query samples. After that, we randomly sample the artificial preferences from the simplex $\Delta=\{\boldsymbol{\lambda}\in\mathbb{R}_{+}^{m}\mid\|\boldsymbol{\lambda}\|_{1}=1\}$ to derive the aggregation targets $g$ given query inputs and preferences. The aggregation targets are then masked to formulate a supervised prediction problem, enabling the model to learn to predict the preference-wise aggregated posterior distribution $q_{\boldsymbol{\theta}}(g|\boldsymbol{x},D_{n};\boldsymbol{\lambda})$. We continuously synthesize data through extensive sampling and optimize the model parameters to minimize the cross-entropy loss function defined in Equation (5), which is equivalent to the expected KL divergence between the true posterior distribution and its approximation, up to an additive constant (Müller et al. 2021). The detailed information and hyperparameter settings regarding the foundational model and the pre-training process can be found in Table 2.
C.2 Visualization of Aggregated Posteriors
Figure 5: Details of the normalized objective functions for the ToyRobust problem, the aggregated posteriors (green solid lines with shaded regions) conditioned on 5 observations and 5 preferences, and the corresponding UCB acquisition functions (red solid lines).
For the convenience of visualization, we use a simple one-dimensional ToyRobust problem (Daulton et al. 2022) as an example to demonstrate that the pre-trained foundation model can effectively learn to infer the aggregated posteriors on unknown problems, as illustrated in Figure 5. The results show that the foundation model provides reliable predictions given the evaluated trajectory of the out-of-distribution problem without any model training and update, allows the model to generalize beyond the synthetic training data. The subproblems defined by adjacent preferences exhibit similar characteristics in terms of the scalarization functions, the predicted mean captures this similarity effectively. Furthermore, the prediction provides a reliable measure of calibration uncertainty, enabling the acquisition function such as upper confidence bound (UCB) shown in the figure to effectively balance exploration and exploitation during the search for candidate solutions.
C.3 Derivation of UHVI Acquisition Function
@@
The derivation mainly depends on two types of definitions on Tchebycheff scalarizations and the theorem on hypervolume scalarization. We consider the the following multi-objective optimization problem (MOP):
@@
where $\boldsymbol{\lambda}=(\lambda_{1},\lambda_{2},\ldots,\lambda_{m})^{T}$ represents a preference sampled from the simplex $\Delta=\{\boldsymbol{\lambda}\in\mathbb{R}_{+}^{m}\mid\|\boldsymbol{\lambda}\|_{1}=1\}$, we can identify each Pareto optimal solution by minimizing $s_{\boldsymbol{\lambda}}^{\text{tch}}(\boldsymbol{x})$ with a specific but unknown preference (Choo and Atkins 1983).
@@
Theorem 1 (Hypervolume Scalarization).
Let $D_{n}=\{\boldsymbol{x}_{i},\boldsymbol{y}_{i}\}_{i=1}^{n}$ be a set of $n$ evaluated samples, $\boldsymbol{z}^{*}=(z_{1}^{*},\cdots,z_{m}^{*})^{T}$ is the ideal point, $\boldsymbol{r}=(r_{1},\cdots,r_{m})^{T}$ is a pre-defined reference point. The hypervolume of the evaluated samples can be expressed as an expectation of Tchebycheff scalarization function over a set of uniformly sampled reference vectors $S=\{\boldsymbol{w}\in\mathbb{R}_{+}^{m}\mid\|\boldsymbol{w}\|_{2}=1\}$ (Deng and Zhang 2019; Zhang and Golovin 2020; Zhang et al. 2023):
@@
Lemma 1 (Hypervolume Improvment with Uncertainty).
Given the the distribution of the aggregation target $g_{\boldsymbol{\lambda}}(\cdot)=-s_{\boldsymbol{\lambda}}^{\textnormal{tch}}(\cdot)$ conditioned on evaluated samples $D_{n}=\{\boldsymbol{x}_{i},\boldsymbol{y}_{i}\}_{i=1}^{n}$, the utility of hypervolume improvement in the normalized objective space at any query point $\boldsymbol{x}$ can be estimated by the following expectation:
@@
Intuitively, maximizing $\alpha_{\textnormal{UHVI}}$ utility is equivalent to maximizing the predicted upper bounds of aggregation targets across all preference simultaneously. We find that our utility exhibits similar expression to that presented in (Song et al. 2024b). However, the key difference is that the formulation of (Song et al. 2024b) is based on the Gaussian process (GP) posterior for each objective, whereas our acquisition function is directly derived from infinite preference-wise aggregated posteriors, which is intractable for traditional GP training and updates.
Appendix D Experiment Setting Details
D.1 Details of Benchmark
Table 3: Detailed information for synthetic problems and real-world applications.
Name $d$ $m$ Type Pareto Front Shape
ZDT1 8 2 Continuous Convex
ZDT2 8 2 Continuous Concave
ZDT3 8 2 Continuous Disconnected
ZDT4 8 2 Continuous Convex
ZDT6 8 2 Continuous Concave
Omnitest 2 2 Continuous Convex
RE21 (Four bar truss design) 4 2 Continuous Convex
RE22 (Reinforced concrete beam design) 3 2 Mixed Mixed
RE23 (Pressure vessel design) 4 2 Mixed Mixed, Disconnected
RE24 (Hatch cover design) 2 2 Continuous Convex
RE25 (Coil compression spring design) 3 2 Mixed Mixed, Disconnected
RE31 (Two bar truss design) 3 3 Continuous Unknown
RE32 (Welded beam design) 4 3 Continuous Unknown
RE33 (Disc brake design) 4 3 Continuous Unknown
RE34 (Vehicle crashworthiness design) 5 3 Continuous Unknown
RE35 (Speed reducer design) 7 3 Mixed Unknown
RE36 (Gear train design) 4 3 Integer Concave, Disconnected
RE37 (Rocket injector design) 4 3 Continuous Unknown
We firstly utilize synthetic problems to evaluate the effectiveness of our method. The tested synthetic problems include ZDT1-4, ZDT6 (Zitzler, Deb, and Thiele 2000) and Omnitest (Deb and Tiwari 2008) test suits. We also conduct experiments on 12 real-world applications from the popular RE test suit (Tanabe and Ishibuchi 2020), which consists multi-objective engineering design problems from various fields with different characteristics. The detailed information can be referred to Table 3.
D.2 Details of Implementation
In many real-world scenarios, the ranges of each objective can vary significantly, sometimes even by orders of magnitude, making Gaussian Process modeling and subsequent optimization processes challenging. To this end, before executing each algorithm, we normalize each $i$-th objective observations $y_{i}(\boldsymbol{x})$ based on the evaluated samples $D_{n}=\{(\boldsymbol{x}_{i},\boldsymbol{y}_{i})\}_{i=1}^{n}$ using $(y_{i}(\boldsymbol{x})-y_{i}^{\text{min}})/(y_{i}^{\text{max}}-y_{i}^{\text{min}})$, where $y_{i}^{\text{min}}=\min_{\boldsymbol{x}\in\mathcal{D}_{n}}y_{i}(\boldsymbol{x})$ and $y_{i}^{\text{max}}=\max_{\boldsymbol{x}\in\mathcal{D}_{n}}y_{i}(\boldsymbol{x})$. When assessing the quality of the candidate solutions obtained after the optimization is completed, we use the true objective observations to calculate the IGD+ or normalized hypervolume metrics, as detailed in Section D.3.
D.3 Details of Performance Metrics
Since the tested synthetic functions have clear analytical expressions with known Pareto fronts, we use the inverted generational distance plus (IGD+) indicator (Ishibuchi et al. 2015) as the metric to evaluate the quality of the final solutions obtained by each algorithm. Let $\mathcal{P}$ be the approximated Pareto front obtained by the tested algorithm, and $\mathcal{P}^{*}$ be a set of well-distributed objective vectors sampled from the true Pareto front. The definition of IGD+ is as follows, indicating a smaller IGD+ value represents the better performance of an algorithm:
$\operatorname{IGD}^{+}(\mathcal{P}^{*},\mathcal{P})=\frac{1}{|\mathcal{P}^{*}|}\sum_{\boldsymbol{v}\in\mathcal{P}^{*}}\min_{\boldsymbol{u}\in\mathcal{P}}\sqrt{\sum_{i=1}^{m}([u_{i}-v_{i}]_{+})^{2}}.$ (21)

### v2

@@ 5.2 Parallel Scalability
In addition to sequential optimization, we further investigate the scalability of our method in the context of parallel optimization. For preference-based acquisition functions, we generate a batch of $q$ candidates by conditioning $q$ random preferences sampled in each iteration. For preference-free variants, since approximating HV or R2 improvement via a finite set of random preferences introduces inherent stochastic deviation, we leverage this characteristic by independently repeating UHVI or UR2I calculations $q$ times to produce $q$ candidates for evaluation. We evaluate this parallel scheme under batch sizes of $q=5$ and $q=10$. Our results, detailed in Appendix D.2, demonstrate that even with these straightforward parallelization strategies, our preference-based methods consistently outperform MOBO algorithms that rely on well-designed batch selection heuristics. This performance gap highlights the advantage of using a foundation model pre-trained on a large variety of synthetic landscapes for optimization. Moreover, it demonstrates that modeling the preference-wise aggregated posterior is inherently well-suited for parallelization, providing a more robust and natural framework for multi-objective optimization without sophisticated batch selection processes.
5.3 Runtime Analysis
Table 4: Comparison of average query time per iteration for all algorithms in sequential optimization ($q=1$). The best, second-best, and third-best mean results for each case are highlighted with dark gray, gray, and light gray backgrounds, respectively.
Algorithm 2 Objectives 3 Objectives 4 Objectives 6 Objectives
MOEA/D-EGO 38.6409 46.8428 49.9341 54.1983
TSEMO 5.6616 6.1670 5.8793 6.3524
USEMO-EI 8.1582 6.6809 7.2214 6.7517
DGEMO 36.2580 95.3770 N/A N/A
qParEGO 11.9262 19.0253 85.5916 185.4954
qNEHVI 10.1029 22.0746 62.9022 78.2812
JES 10.3307 33.4365 41.6235 44.9701
BOFormer 0.6679 0.7483 N/A N/A
FoMEMO-EI 2.5706 3.9841 3.1434 2.8501
FoMEMO-UCB 1.7201 1.1666 1.1155 2.0801
FoMEMO-UHVI 0.6189 0.6094 0.6197 0.7432
FoMEMO-UR2I 1.6248 1.7077 1.3686 1.1716
We evaluate the computational efficiency in sequential optimization across 4 engineering design problems (four bar truss, speed reducer, car side impact design and water resource planning problems) with objective counts ranging from 2 to 6, as shown in Table 4. The reported query time is averaged from 10 independent runs with 100 iterations each. Beyond the inherent time advantage of bypassing model retraining, here we ensure a rigorous comparison by monitoring only the duration required for acquisition function optimization to generate candidates, excluding the time allocated to model retraining (where applicable) and true function evaluations. This allows us to isolate and highlight the efficiency gains provided by the foundation model’s in-context optimization capabilities.
Due to the inference overhead of GP models, the computational cost of associated operations—such as the population-based optimization in MOEA/D-EGO, the intricate batch selection in DGEMO, and the expensive HV calculations in qNEHVI—increases drastically with problem complexity. In contrast, benefiting from the fast in-context learning of our foundation model, FoMEMO demonstrates superior time efficiency in optimizing acquisition functions as the objective dimension increases, some even achieving speedups of several orders of magnitude in high-dimensional scenarios. While BOFormer remains competitive in 2D and 3D objectives, it is still surpassed by FoMEMO-UHVI. Although the true function evaluations often dominate the model inference time in real-world problems, these results highlight the substantial efficiency gains provided by the foundation model paradigm, offering a more robust and scalable framework for many real-world challenges.
5.4 Modeling Accuracy Analysis
Since the efficacy of our method relies on modeling fidelity, in addition to providing a posterior visualization of a simple problem (see Appendix A.2), we further conduct a statistical analysis of the model’s ability to fit aggregation targets across problems with varying objective dimensions. Specifically, we monitor the accuracy of the predicted means throughout the training process by partitioning 1000 randomly generated datasets into training and testing sets using a 7:3 ratio. For each of the five independent experimental trials, a preference vector is randomly sampled to generate the aggregation targets. We employe RMSE and MAE metrics to quantify point-wise accuracy, alongside the $R^{2}$ score to evaluate the model’s capacity to capture the underlying data trends. We observe that as training progresses, the model’s predictive accuracy on unseen problems improves consistently. This evolution in modeling fidelity provides a fundamental explanation for the superior optimization performance demonstrated by our method.
Figure 2: The evolution of modeling accuracy in terms of regression error metrics (RMSE, MAE) and $R^{2}$ score during the training process.
@@ 6 Conclusion and Future Work
We propose a novel paradigm named FoMEMO, which, to the best of our knowledge, represents the first attempt to establish a foundation model to address a wide range of expensive MOPs in an in-context manner. Extensive experiments verify that our method exhibits remarkable adaptability and generalization in various synthetic problems and real-world applications, showing superior performance and high efficiency in most cases compared to existing methods.
@@
In the future, we expect that this work opens up opportunities for many exciting directions worth exploring: (i) Develop architectures that allow for datasets with larger sizes, including embedding techniques and training schemes that efficiently handle higher-dimensional feature and objective spaces. (ii) Develop efficient acquisition functions tailored for high-dimensional and high-throughput parallel scenarios. (iii) Investigate whether fine-tuning can help improve the performance in domain-specific tasks.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Model and Algorithm Details
A.1 Details of Foundation Model and Pre-training
@@
Maximum Sample Length 256
Model Training Batch Size 128
@@
Learning Rate $5\times 10^{-5}$
Our foundational model mainly employs the recently proposed Prior-Data Fitted Networks (PFNs) (Müller et al., 2021) as the basic architecture, which is based on a Transformer encoder (Vaswani et al., 2017). We utilize linear layers to encode trajectory and query points, along with sampled preferences, as contextual inputs for the foundation model. Following the methods outlined in (Hollmann et al., 2022) and (Müller et al., 2023), we set a maximum dimension $K$ for encoding the inputs, which are applicable to both feature and objective dimensions. We zero-pad the sampled dimension $k$ and linearly scale the dimension by a factor of $K/k$ to ensure that the magnitudes of the input encodings are consistent across different dimensions. Since different query points should not depend on one another, an attention mask is implemented to allow each trajectory point can only attend to each other, while the query points can only attend to the trajectory points. To maintain the permutation invariance of the input dataset, the positional encodings are removed and the input points are fed to the model as the sum of their encoded embeddings. During training, we maintain a maximum sample length of $N$ with $n$ trajectory points and NnN-n query points, as used in (Müller et al., 2021). We sample each $n$ from $\{1,...,N-1\}$ with the weight of 1/(Nn)1/(N-n) to simulate varying trajectory sizes during the optimization process, while ensuring that each trajectory size has approximately the same number of query points for training. To model the continuous distribution for prediction, the regression head of PFNs discretizes the continuous distribution into a piecewise constant distribution, referred to as the Riemann distribution (Müller et al., 2021). This helps transform the model training from a regression task to a classification problem. The boundaries of each discrete interval are estimated by sampling a large synthetic dataset from the same synthetic prior, ensuring that each interval maintains equal probability in the prior data. Additionally, the last bar on each side is replaced with a suitably scaled half-normal distribution to enhance training stability. To account for the distributional shifts in aggregated posteriors across varying objective dimensions, we propose an objective-aware regression head. This architecture dynamically routes latent representations to specialized regressor conditioned on the specific objective dimension of the problem. By leveraging this dimension-specific mapping, the model can more accurately capture the differences of the aggregated posteriors with varying objective dimensions, thereby significantly bolstering predictive precision when generalizing to unseen optimization landscapes across diverse objective dimensions.
@@
where $\sigma$ and $l$ represent the output scale and length scale kernel parameters, respectively. We also sample the length scale parameter for each feature dimension during the latent function sampling, along with an random observation noise to enhance function diversity. In the third step, we randomly sample input features from the $d$-dimensional unit cube, and then evaluate the sampled latent functions to generate the corresponding observations, resulting in $n$ trajectory samples and NnN-n query samples. After that, we randomly sample the artificial preferences from the simplex $\Delta=\{\boldsymbol{\lambda}\in\mathbb{R}_{+}^{m}\mid\|\boldsymbol{\lambda}\|_{1}=1\}$ to derive the aggregation targets $g$ given query inputs and preferences. The aggregation targets are then masked to formulate a supervised prediction problem, enabling the model to learn to predict the preference-wise aggregated posterior distribution $q_{\boldsymbol{\theta}}(g|\boldsymbol{x},D_{n};\boldsymbol{\lambda})$. We continuously synthesize data through extensive sampling and optimize the model parameters to minimize the cross-entropy loss function defined in Equation (5), which is equivalent to the expected KL divergence between the true posterior distribution and its approximation, up to an additive constant (Müller et al., 2021). The detailed information and hyperparameter settings regarding the foundational model and the pre-training process can be found in Table 5.
A.2 Visualization of Aggregated Posteriors
Figure 3: Details of the normalized objective functions for the ToyRobust problem, the aggregated posteriors (green solid lines with shaded regions) conditioned on 7 observations and 5 preferences, and the corresponding UCB acquisition functions (red solid lines).
To facilitate visualization, we employ a one-dimensional ToyRobust problem (Daulton et al., 2022) as a representative case to demonstrate the foundation model’s ability to predict aggregated posteriors with unseen objective functions, as illustrated in Figure 3. The results indicate that the model provides reliable posterior estimates directly from the evaluated trajectory of the unknown problem, without requiring further training or parameter updates. This underscores the model’s strong generalization capabilities beyond its training distribution. Specifically, the predicted mean accurately captures the global trends of individual sub-problems and also maintains similar local characteristics defined by adjacent preferences. Furthermore, the model yields well-calibrated uncertainty estimates, enabling acquisition functions—such as the Upper Confidence Bound (UCB) shown in the figure—to effectively balance exploration and exploitation throughout the optimization process.
A.3 Derivation of UHVI Acquisition Function
@@
The derivation mainly depends on two types of definitions on Tchebycheff scalarizations and the theorem on hypervolume scalarization. We consider the following multi-objective optimization problem (MOP):
@@
where $\boldsymbol{\lambda}=(\lambda_{1},\lambda_{2},\ldots,\lambda_{m})^{T}$ represents a preference sampled from the simplex $\Delta=\{\boldsymbol{\lambda}\in\mathbb{R}_{+}^{m}\mid\|\boldsymbol{\lambda}\|_{1}=1\}$, we can identify each Pareto optimal solution by minimizing $s_{\boldsymbol{\lambda}}^{\text{tch}}(\boldsymbol{x})$ with a specific but unknown preference (Choo and Atkins, 1983).
@@
Theorem A.1 (Hypervolume Scalarization).
Let $D_{n}=\{\boldsymbol{x}_{i},\boldsymbol{y}_{i}\}_{i=1}^{n}$ be a set of $n$ evaluated samples, $\boldsymbol{z}^{*}=(z_{1}^{*},\cdots,z_{m}^{*})^{T}$ is the ideal point, $\boldsymbol{r}=(r_{1},\cdots,r_{m})^{T}$ is a pre-defined reference point. The hypervolume of the evaluated samples can be expressed as an expectation of Tchebycheff scalarization function over a set of uniformly sampled reference vectors $S=\{\boldsymbol{w}\in\mathbb{R}_{+}^{m}\mid\|\boldsymbol{w}\|_{2}=1\}$ (Deng and Zhang, 2019; Zhang and Golovin, 2020; Zhang et al., 2023):
@@
Lemma A.2 (Uncertainty-aware Hypervolume Improvement).
Given the distribution of the aggregation target $g_{\boldsymbol{\lambda}}(\cdot)=-s_{\boldsymbol{\lambda}}^{\textnormal{tch}}(\cdot)$ conditioned on evaluated samples $D_{n}=\{\boldsymbol{x}_{i},\boldsymbol{y}_{i}\}_{i=1}^{n}$, the utility of hypervolume improvement in the normalized objective space at any query point $\boldsymbol{x}$ can be estimated by the following expectation:
@@
Intuitively, maximizing $\alpha_{\textnormal{UHVI}}$ utility is equivalent to maximizing the predicted upper bounds of aggregation targets across all preference simultaneously. We find that our utility exhibits similar expression to that presented in (Song et al., 2024b). However, the key difference is that the formulation of (Song et al., 2024b) is based on the Gaussian process (GP) posterior for each objective, whereas our acquisition function is directly derived from the infinite preference-wise aggregated posteriors, which is intractable for traditional GP training and updates.
A.4 Derivation of UR2I Acquisition Function
Theorem A.3 (R2 Indicator).
Let $D_{n}=\{\boldsymbol{x}_{i},\boldsymbol{y}_{i}\}_{i=1}^{n}$ be a set of $n$ evaluated samples, the R2 indicator of the evaluated samples is defined as an expectation of Tchebycheff scalarization function over a set of preferences sampled from the simplex $\Delta=\{\boldsymbol{\lambda}\in\mathbb{R}_{+}^{m}\mid\|\boldsymbol{\lambda}\|_{1}=1\}$ (Hansen and Jaszkiewicz, 1994; Brockhoff et al., 2012; Zhao et al., 2024):
$\textnormal{R}_{2}=\mathbb{E}_{\boldsymbol{\lambda}\sim\Delta}[\min_{1\leq i\leq n}s_{\boldsymbol{\lambda}}^{\textnormal{tch}}(\boldsymbol{x}_{i})].$ (22)
Analogous to the derivation of the UHVI acquisition function presented above, the utility for the UR2I acquisition function can be naturally derived as follows. For the sake of brevity, the formal proof is omitted here.
Lemma A.4 (Uncertainty-aware R2 Improvement).
Given the distribution of the aggregation target $g_{\boldsymbol{\lambda}}(\cdot)=-s_{\boldsymbol{\lambda}}^{\textnormal{tch}}(\cdot)$ conditioned on evaluated samples $D_{n}=\{\boldsymbol{x}_{i},\boldsymbol{y}_{i}\}_{i=1}^{n}$, the utility of R2 improvement in the normalized objective space at any query point $\boldsymbol{x}$ can be estimated by the following expectation:
$\alpha_{\textnormal{UR2I}}(\boldsymbol{x})=\mathbb{E}_{\boldsymbol{\lambda}\sim\Delta}[\max\{0,\min_{1\leq i\leq n}\{-g_{\boldsymbol{\lambda}}(\boldsymbol{x}_{i})\}-[-\textnormal{UCB}(g_{\boldsymbol{\lambda}}(\boldsymbol{x}))]\}].$ (23)
Appendix B Experimental Setup Details
B.1 Benchmark Problems
Table 6: Detailed information of test benchmarks.
Type Name Feature Dimension ($d$) Objective Dimension ($m$)
Synthetic Function BC (Branin-Currin) 2 2
AR (Ackley-Rosenbrock) 2 2
ABC (Ackley-Branin-Currin) 2 3
ABD (Ackley-Branin-Dixon) 2 3
ACD (Ackley-Currin-Dixon) 2 3
BCD (Branin-Currin-Dixon) 2 3
Engineering Design Four Bar Truss Design 4 2
Pressure Vessel Design 4 2
Hatch Cover Design 2 2
Disc Brake Design 4 3
Speed Reducer Design 7 3
Gear Train Design 4 3
Rocket Injector Design 4 3
Car Side Impact Design 7 4
Conceptual Marine Design 6 4
Water Resource Planning 3 6
Hyperparameter Optimization Lego 5 2
Materials 5 2
Mic 5 3
ship 5 3
To evaluate the performance and generalization of our method, we employ a diverse suite of synthetic and real-world benchmarks, including 6 synthetic functions, 10 real-world engineering design problems and 4 hyperparameter optimization tasks. Specifically, we firstly utilize a combination of synthetic black-box functions characterized by a wide spectrum of landscapes, including unimodal versus multimodal structures, smooth versus non-smooth surfaces, and challenges involving narrow valleys and complex variable coupling, we utilize the implementation in BOFormer (Hung et al., 2025). Secondly, we incorporate the RE benchmark suite (Tanabe and Ishibuchi, 2020), which comprises real-world multi-objective engineering design problems from various fields such as pressure vessel, speed reducer, and rocket injector design, covering multi-objective ($m$ = 2, 3) to many-objective ($m$ = 4, 6) scenarios. Detailed description and characterization of the objective functions for these problems can be found in (Tanabe and Ishibuchi, 2020). Finally, we validate our approach on a hyperparameter optimization (HPO) problem for 3D Gaussian Splatting (3DGS). This problem aims to identify the optimal configurations of the the recent state-of-the-art method 3DGS (Kerbl et al., 2023) in 3D object reconstruction task (Mildenhall et al., 2021). The optimization objectives encompass rendering quality (measured by PSNR), model size, and the total number of Gaussian primitives. To comprehensively assess performance, we conduct evaluations across four diverse reconstruction scenes: Lego, Materials, Mic, and Ship from (Mildenhall et al., 2021). The detailed information of the benchmarks can be referred to Table 6.
B.2 Algorithm Hyperparameters
Since qNEHVI, qParEGO, JES, and our proposed method are all implemented based on the BoTorch framework (Balandat et al., 2020), we maintain consistent optimization settings for their respective acquisition functions. Specifically, we employ the multi-start L-BFGS-B optimizer with 20 restarts seeded from 1024 pseudo-random samples through BoTorch’s initialization heuristic, consistent with the protocol in (Ament et al., 2023). To strike a balance between optimization accuracy and computational efficiency, we utilize a stochastic approximation for the UHVI and UR2I utilities by randomly sampling 100 preference vectors during each optimization step. For all other baselines, we adhere to the default parameter configurations provided in their original code implementations unless otherwise specified.
B.3 Performance Metrics